OpenStudy (anonymous):

What is the general form of a polynomial p(x) of degree 2m such that p(x^2)=p(x)p(x+1)

OpenStudy (anonymous):

\[P(x)=(x^2-x+1)^m\]

OpenStudy (anonymous):

Are you sure?

OpenStudy (anonymous):

wait a sec

OpenStudy (anonymous):

this time i am thinking that \(p(0)\) unless this is a constant function

OpenStudy (anonymous):

@mukushla \[ \left\{p\left(x^2\right)=x^4-x^2+1\\p(x) p(x+1)=x^4+x^2+1\right\} \] in your solution for m=1

OpenStudy (anonymous):

since as before since \(p(0)=p(0)p(1)\implies p(0)=0\) or \(p(1)=1\) but if \(p(1)=1\) then since \(p(1)=p(1)p(2)\) we have \(p(1)=0\) which it is not or \(p(2)=p(1)\)

OpenStudy (anonymous):

The only constant solutions are p(x) = 0 and p(x) = 1 for all x. So we restrict consideration to the case where p(x) is nonconstant. now for x = 0 and x = −1 the equation becomes p(0) = p(0) p(1), p(1) = p(−1) p(0) . then 0 is a root if and only if 1 is a root. so i think P(x) is in th form \[P(x)=a x^l (1-x)^k Q(x)\]

OpenStudy (anonymous):

but we have to find Q(x) !!!

OpenStudy (anonymous):

zero is a root yes? which makes 1 a root as well

OpenStudy (anonymous):

let x=-1 p(1) = p(−1) p(0)

OpenStudy (anonymous):

right, and since \(p(0)=0\) 1 is a root also

OpenStudy (anonymous):

unless my logic was faulty above

OpenStudy (anonymous):

yes thats right

OpenStudy (anonymous):

now i think we can put the P(x) in the equation to get the Q(x)

OpenStudy (anonymous):

let me do that

OpenStudy (anonymous):

coefficient of largest exponent of x must be equal on 2 sides so a=a^2 then a=1

OpenStudy (anonymous):

\[ax^{2l} (x − 1)^l (x + 1)^l Q(x^2) = a^2 x^{l+k} (x + 1)^l (x − 1)^k Q(x) Q(x + 1)\] i got rid of the large so i could read it

OpenStudy (anonymous):

any way to use that the sum of the coefficients is zero?

OpenStudy (anonymous):

also as your argument before, if \(r\) is a root, then so is \(r^2\) limiting the possible zeros

OpenStudy (anonymous):

i dont know how to use that (sum of the ...=0)

OpenStudy (anonymous):

it seems l=k but i cant prove it

OpenStudy (anonymous):

looks similar to previous one only possible roots are 0, 1, \(-\frac{1}{2}\pm\frac{\sqrt{3}}{2}i\) by what you wrote in previous post that is, if \(r\) is a root, so is \(r^2, r^4, ...\)

OpenStudy (anonymous):

the problem is 0,1 in the previous one 0,1 were not roots

OpenStudy (anonymous):

i guess i meant only other possible roots

OpenStudy (anonymous):

maybe these are the only two which is what i suspect a quick calculation tells me for example that \(p(x)=x(1-x)\) works whereas \[p(x)=x(1-x)(x^2+x+1)=x-x^4\] does not

OpenStudy (anonymous):

oh yes sorry u mean to find all possible roots for this one

OpenStudy (anonymous):

i am off to the movies love to know what the solution is, so i'll check back later

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok i want to prove 0,1 are only roots of such polynomials similar to previous one Suppose p(a) = 0 for some complex number a. Then p(a^2) = p(a^4) =· · · = 0. then |a| = 1. Similarly, if a is any such root, then P((a − 1)^2) = P(a − 1) P(a) = 0. So (a − 1)^2 is also a root. It follows that |a − 1| = 1. there are only 2 complex numbers such that |a| = 1 and |a − 1| = 1 . \[a=\frac{1}{2}\pm \frac{\sqrt{3}}{2} i\] but unlike previous one \[a^2=-\frac{1}{2}\pm \frac{\sqrt{3}}{2} i \neq a\] that means only ossible roots are 0 and 1 so \[P(x)=c x^l (x-1)^k\] put this in the equation we get \[cx^{2l}(x−1)^k(x+1)^k=c^2x^{l+k}(x+1)^l(x−1)^k \ \ \ \ \ \ (I)\] leading coefficient of 2 sides must be equal then c=c^2 ----> c=1 from (I) \[(x(x+1)(x-1))^{l-k}=1 \rightarrow \rightarrow l=k=m\] so the only answer is \[\large P(x)=x^m (x-1)^m\]

OpenStudy (anonymous):

@eliassaab @satellite73

OpenStudy (anonymous):

@mukushla Excellent solution.

OpenStudy (anonymous):

tnx